OpenStudy (anonymous):

Janette went to Tire Mart and paid $122.57 per tire for 4 tires and a flat rate fee for installation for a total of $555.28 before taxes. Monique went to Fast Wheels and paid $136.71 per tire for 4 tires and a flat rate fee for installation for a total of $585.34 before taxes. The difference in the amount that they paid in installation charges is $ .

OpenStudy (texaschic101):

Janette : 122.57 for 4 tires = (122.57 * 4) = 490.28 490.28 + x (your flat rate fee) = 555.28 x = 555.28 - 490.28 x = 65 Monique : 136.71 for 4 tires = (136.71 * 4) = 546.84 546.84 + x (your flat rate fee) = 585.34 x = 585.34 - 546.84 x = 38.50 so the difference is : 65 - 38.50 = $26.50 questions ?

OpenStudy (anonymous):

Nope thank you

OpenStudy (texaschic101):

no problem :)